Two mercury drops (each of radius ‘r’ ) merge to form bigger drop. The surface energy of the bigger drop, if T is the surface tension, is:
(a) 4pr2T
(b) 2pr2T
(c) 28/3pr2T
(d) 25/3pr2T

Grade:11

1 Answers

Kevin Nash
9 years ago

(c) Sum of volumes of 2 smaller drops = Volume of the bigger drop

2.4/3pr3 = 4/3pR3 ? R = 21/3r

Surface energy = T.4pR2

= T4p2 2 = T.28/3 pr2